# Building a continuous benchmarking ecosystem in bioinformatics

### v1

@@ Introduction
Software plays an important role in the current scientific landscape [30], since it is widely used throughout the scientific process, including data collection, simulation, analysis and reporting. Researchers are encouraged to publish their data and code to enhance transparency and reproducibility in their work. Ideally, this would increase benchmark (software) reuse by the wider scientific community. In practice, however, it is sometimes easier to develop new code than to reuse existing software [31]. This leads to the phenomenon of ‘academic abandonware’, where projects are forgotten in code repositories, exacerbating the reproducibility crisis in research; one of the noted challenges in benchmarking was the lack of extensible frameworks [2]. A common reason for abandoned projects is their failure to adhere to the FAIR (Findable, Accessible, Interoperable, Reusable) principles [6]. Another common cause is the life cycle inherent to academic practice: after delivering a research output, there are often no resources to maintain a software package in the longer term; left alone, the likelihood of a particular library compiling or being able to run after 5-10 years will drastically decrease.
Operational costs are given by the expected compute and storage requirements. In the context of benchmarking, predicting CPU and memory usage is challenging. The methods under evaluation typically have variable computational demands. Benchmark designers could impose resource limits, thereby evaluating the methods under different constraints. Storage costs, on the other hand, are more predictable and can be directly influenced by design decisions of the benchmark (e.g., low retention for re-generatable artifacts, ‘cold storage’ for code archives and software, external data stores such as Zenodo for method performance artifacts, CernVM-FS for reusable software).
@@
A good benchmarking system needs to package its components in a way that ensures that their execution and artifacts can be consistently replicated within practical limits.
The UNIX design philosophy [32] is worth embracing to produce software for benchmarking systems that are lightweight, robust, and easily maintainable; ease of use and extensibility are also desirable. Furthermore, to reduce the entry barrier for contributors, the benchmarking system should be capable of running locally with minimal setup. This approach ensures that the system remains relevant and functional even in the absence of centralization. Furthermore, intuitive APIs and sensible defaults [33] are factors that contribute to highly usable systems. Structured approaches to documentation, including adopting frameworks like diataxis [34], can facilitate adoption and new contributions to a benchmark.
@@
Software reproducibility comes from controlling the triad of data, code, and environment [35]. Computationally, a benchmark executes code that transforms input data into outputs: this execution takes place and is affected by the execution environment; the environment encompasses the base system (OS, compiler toolchains, libraries) and a set of configurations. The benchmark definition should control as much as possible of the execution environment.
@@
    The benchmarking system itself has an impact on operational costs, such as storage needs, execution platform, etc. For example, it imposes a certain software stack and base dependencies, and orchestrates the execution of a benchmark plan; usually, the system mandates the choice of a particular workflow manager. At this level, any requirements for input/output formats and shapes are defined; different systems can impose different constraints.
@@
    Contributions related to individual benchmark components (e.g., datasets, methods, metrics) are expected to be small, typically short scripts that process data and wrap functionality by importing external libraries (where the component implementation itself is developed). Even so, imposing good practices at this level (output validation, testing for abnormal termination, ability to run with a subset of input data) can be beneficial to increase the quality and maintainability of the contributions. The toolset can also enforce the validation of metadata annotations (authorship, versioning, licences) for each contributed module.
@@
    The software dependencies have easily the biggest impact on replicability and maintainability; archiving all combinations of a large dependency tree, across an arbitrary number of base OS images, will exponentially increase retention and maintenance costs. Turning the long-term replicability of a benchmark into a tractable problem is linked to choosing a suitable software management system. Several software management systems have emerged to address these problems, from containerization (e.g., apptainer or singularity) [36] and automation of reproducible and efficient software management (EasyBuild [37] or Spack [38]) to distributing software installations such as CernVM-FS [39], or combinations of the two such as computecanada [40] and EESSI [41].
@@
Even the simplest benchmark designed and run by a single person aiming to compare their method against various ‘gold standard’ methods incorporates the work of many (dataset generators, method developers, operating systems with system dependencies and their developers). Benchmark artifacts are generated by running a workflow [42]. Hence, data, code and operating system dependencies are all accessible during the benchmarking process. For reproducibility and transparency and to build trust [43], this freedom should be extended for others to run (re-run), copy (including fork), distribute (e.g., snapshot), study (e.g., read how another person is implementing my own method), change (bug fix, contribute) and improve (extend) benchmarks. Such freedoms are granted (or restricted) by software and content licenses, as are copyright and authorship and attribution [44]. Without explicit licenses, even publicly available code is copyrighted by default, and reuse is restricted (to reproduce, distribute, or create derivative works) [44].
In practical terms, a benchmarking system needs to be aware of software and data licenses, track further changes and attribute them to the (academic) benchmark contributor. A concise summary of license rights and restrictions, and compatibility across commonly-used open-source licenses can be found in Table I. Benchmark contributions should bring a clear FOSS (free and open source software) license, perhaps by assigning the copyright holder to contributors describing their name, email and ORCID (as the canonical scholarly identifier), hence allowing seamless reuse, transparency and easing attribution.
Table I: Overview of commonly used open-source licenses in benchmarking studies. Code or data with restrictions to redistribute or modify are not suitable for collaborative benchmarking. License: license name. Redistribution: right to redistribute, e.g. reuse in another benchmark/somewhere else. Modification: right to modify the code to produce derivative works, e.g. in another benchmark/somewhere else. Sublicensing: for derivative works, obligation to keep the same license or freedom to be licensed differently (even copyright). Relevant if the code contributor does not want derivative works to be ‘closed source’ or copyrighted. Commercial use: for derivative works, whether commercial usage is restricted or not. Relevant to code contributors not willing commercial reuse of their code (by others). Attribution: for derivative works, obligation to attribute original author(s). Relevant to code contributors willing their names/authorship to be credited in derivative works. Linking: dynamic or static linking / use as libraries. Particularly important for metrics, as these code can be frequently incorporated (=executed) across benchmarks.
@@
Similarly, even if all benchmark code were free software, running it could be hindered by (non-free) proprietary interpreters (e.g., MATLAB). Benchmarking systems should be explicit in allowing or disallowing proprietary software requirements, and be aware and track EULAs (end user license agreements). This conundrum extends to the operating system layer and to system dependencies (e.g., compilers). In terms of benchmark re-runnability, fully FOSS system dependencies arguably offer the highest reusability, and can be handled in a reproducible manner across operating systems and system architectures (e.g., amd64, arm64) [41].
Finally, data (inputs and outputs) are frequently stored in centralized data repositories [45, 46, 47] under content licenses, which might enforce attribution or copyleft. A benchmarking system should automate license metadata handling during data import and export, crediting the original authors if applicable; it also represents an opportunity to enforce that academics use appropriate licenses for data and code.
Scientific benchmarking is rarely a single-person endeavour and has to deal with diverse content and contributions. Software and content licenses standardize the (legally binding) usage rights and restrictions. Benchmarking systems can leverage licenses and version control to track, store and monitor the reuse of data and code in an automated way. On the other hand, contribution attribution in the academic sense (e.g., authorship) is as important and beyond automation. Collaborative benchmarking initiatives should explicitly list guidelines on using and distributing code and results (i.e., licenses) as well as clearly communicate the authorship strategies. These guidelines, complemented with a code of conduct, could be bundled into the benchmark description.
@@
So far, no single benchmarking system can accommodate all use-cases in bioinformatics. However, documentation of design choices for a given system can help to align expectations. Some important design trade-offs are:
  • Flexibility vs. complexity: unrestricted freedom for method contributors can mean displacing the burden of maintainability towards the benchmarker. Fewer restrictions (e.g., simpler entry points) means that more methods can be included in a single benchmark; but, it can also imply a higher complexity for the initial setup by the benchmarker. Constraints can also enable validation, which can greatly help with the development process and quality assurance.
  • Perfect software reproducibility is possible [48], but it comes with a high engineering cost. It is thus reasonable for a benchmarker to state the desired degree of replicability and the corresponding constraints; different tasks might demand different levels of compliance.
  • Security concerns: decentralized runs lower the participation barrier, but it also increases the potential for attacks on their computing environment. Sandboxed environments (restricting network access, filesystem access etc) can be provided as mitigation. Additionally, static and dynamic code analysis in the continuous integration process can aid in approving potentially risky submissions.
@@
Benchmarking involves applying methods to data. The types of data can be diverse and even a single dataset can be stored in multiple formats. A broad distinction can be made according to general vs. specific formats and open vs. proprietary formats. File formats specific to a programming language should be avoided since those formats may change over time, security risks can arise [49, 50], and in general they are not interoperable. Likewise, proprietary formats should be avoided, since they may change and could require specific (proprietary) software to read. General language-agnostic formats often do not change over time and are thus more interoperable such that these datasets can be made FAIR [6]. Generally, using established standards is recommended (e.g., SAM, BED, FASTQ, etc for genomics), since related methods are developed around these (typically storage-efficient) formats.
If raw input data consists of multiple file formats, an option could be to convert the data into a common format with the caveat that this ‘intermediate’ raw data could be either stored (increased storage usage, easier benchmark definition) or created ‘on the fly’ (increased compute). Furthermore, how raw input data is retrieved (e.g., automatic retrieval once and store locally, automatic retrieval every time, or manual retrieval and store locally) needs to be decided.
Additionally, which level and scope of data validation needs to be decided. Validations include content integrity (e.g., md5sum), adherence to a data specification (e.g., JSON Schema), and data semantics (e.g., SHACL for linked data). Intermediate data and results should be made FAIR, leading to additional decisions, including the scope of data sharing (e.g., only the most important subset), the handling of the metadata (automatic versus manual annotation) and the data storage provider and policies (where and how to make available, e.g., GEO or Zenodo). Automatic generation of metadata might require predefined standards and an ontology to fully track how the data was generated [51].
@@ Conclusion
Benchmarking is a cornerstone practice in bioinformatics, impacting how methods are conceived, developed, published and chosen. Benchmarking is driven by the effort of people with the expertise to address (biological and technical) problems. This expertise can be leveraged by current practices of community benchmarking, so scientific advance is eased by joint efforts to catalog open problems, generate (or simulate) appropriate input data, develop methods to address them, and evaluate their performances. Community benchmarking has been proved successful and very productive [52, 12, 13].
Here we discussed the motivation and process of community benchmarking at several levels, from definition to execution, while highlighting technologies to potentially ease the technical, scientific and contribution challenges.

### v2

@@ Introduction
Software plays an important role in the current scientific landscape [33], since it is widely used throughout the scientific process, including data collection, simulation, analysis and reporting. Even if open sourced, however, it is sometimes easier to develop new code than to reuse existing software [34]. This leads to the phenomenon of ‘academic abandonware’, where projects are forgotten in code repositories and not maintained after delivering a research output.
Operational costs of benchmarks are given by the expected compute and storage requirements. In the context of benchmarking, methods under evaluation typically have variable computational demands. Benchmark designers could impose resource limits, thereby evaluating the methods under constraints. Storage costs, on the other hand, are more predictable and can be directly influenced by design decisions of the benchmark (e.g., low retention for re-generatable artifacts, ‘cold storage’ for code archives and software, external data stores such as Zenodo for method performance artifacts, CernVM-FS for reusable software).
@@
A good benchmarking system needs to package its components in a way that ensures that their execution and artifacts can be consistently replicated (within practical limits).
The UNIX design philosophy [35] is worth embracing to produce software for benchmarking systems that are lightweight, robust, and easily maintainable; ease of use and extensibility are also desirable. Furthermore, to reduce the entry barrier for contributors, the benchmarking system should be capable of running locally with minimal setup. This approach ensures that the system remains relevant and functional even in the absence of centralization. Furthermore, intuitive APIs and sensible defaults [36] are factors that contribute to highly usable systems. Structured approaches to documentation, including frameworks such as diataxis [37], can facilitate adoption and new contributions to a benchmark.
@@
Computational reproducibility comes from controlling the triad of data, code, and environment [38]. During benchmarking, a workflow runs code transforming inputs into outputs; execution is affected by the base system (OS, compiler toolchains, libraries) and its configuration. The benchmark definition should control the execution environment.
@@
    The benchmarking system has an impact on operational costs, such as storage needs, execution platform, etc. The system mandates the choice of a particular stack and workflow manager. At this level, any requirements for input/output formats and shapes are defined.
@@
    Benchmark components, like dataset-generating simulations, methods, or metrics, are typically small scripts that process data and use external libraries. Even if method development happens elsewhere, good practices should be embraced (output validation, abnormal termination checks, running on subsets of input data). Metadata validation should be enforced for each module, including authorship, versioning, and licenses.
@@
    Software dependencies cause the biggest impact on replicability and maintainability; archiving all combinations of a large dependency tree, across several OS images, exponentially increases retention and maintenance costs. Making the long-term replicability of a benchmark tractable implies adopting a sane software management system. Several solutions exist, from containerization (e.g., apptainer or singularity) [39] and automation of reproducible and efficient software management (EasyBuild [40] or Spack [41]) to distributing software installations such as CernVM-FS [42], or combinations of the two such as computecanada [43] and EESSI [44].
@@
Even a simple, single-person benchmark comparing methods relies on multiple contributions, from dataset creators and method developers to system software and workflow authors. Hence, data, code and operating system dependencies are all accessible during the benchmarking process. For reproducibility, transparency and to build trust [45], this accessibility and freedom should be extended to others to run (re-run), copy (including fork), distribute (e.g., snapshot), study (e.g., read how someone is implementing a method), change (bug fix, contribute) and improve (extend) benchmarks. Such freedoms are granted (or restricted) by software and content licenses, as are copyright and authorship and attribution [46]. Without explicit licenses, even publicly available code is copyrighted by default, and reuse is restricted (to reproduce, distribute, or create derivative works) [46]. A concise summary of license rights and restrictions, and compatibility across commonly-used open-source licenses can be found in Table 1. Benchmark contributions should be licensed and associated to scholarly names including email and ORCID (as the canonical identifier), hence allowing seamless reuse, transparency and attribution.
Table 1: Overview of commonly used open-source licenses in benchmarking studies. Code or data with restrictions to redistribute or modify are not suitable for collaborative benchmarking. License: license name. Redistribution: right to redistribute, e.g., reuse in another benchmark/somewhere else. Modification: right to modify the code to produce derivative works, e.g., in another benchmark/somewhere else. Sublicensing: for derivative works, obligation to keep the same license or freedom to be licensed differently (even copyright). Relevant if the code contributor does not want derivative works to be ‘closed source’ or copyrighted. Commercial use: for derivative works, whether commercial usage is restricted or not. Relevant to code contributors not willing commercial reuse of their code (by others). Attribution: for derivative works, obligation to attribute original author(s). Relevant to code contributors willing their names/authorship to be credited in derivative works. Linking: dynamic or static linking / use as libraries. Particularly important for metrics, as these code can be frequently incorporated (=executed) across benchmarks.
@@
Similarly, even if all benchmark code were free software, running it could be hindered by (non-free) proprietary interpreters (e.g., MATLAB). Benchmarking systems should be explicit in allowing or disallowing proprietary software, and be aware and track EULAs (end user license agreements). This conundrum extends to the operating system layer and to system dependencies (e.g., compilers). In terms of benchmark re-runnability, fully free software system dependencies arguably offer the highest reusability, and can be handled in a reproducible manner across operating systems and system architectures (e.g., amd64, arm64) [44].
Finally, data files (inputs and outputs) are frequently stored in centralized repositories [47, 48, 49] under content licenses, which might enforce attribution or copyleft. A benchmarking system should automate license metadata handling during data import and export, crediting the original authors if applicable; it also represents an opportunity to enforce that academics use appropriate licenses for data and code.
@@
So far, no single benchmarking system can accommodate all use cases in bioinformatics. However, documentation of design choices for a given system can help to align expectations. Some important design trade-offs are:
  • Flexibility vs. constraints: An unconstrained benchmark grants method contributors greater flexibility but requires more effort from the benchmarker (maintenance, version support). Introducing constraints (such as software dependencies or common data formats) adds initial setup work but can boost maintainability through built-in validation and quality checks.
  • Achieving software reproducibility needs discipline and commitment [50]. A benchmarker can state the desired degree of replicability and constraints.
  • Security concerns: decentralized runs lower the participation barrier, but also increase the attack surface. Sandboxed environments (restricting network access, filesystem access etc) can be provided as mitigation. Additionally, static and dynamic code analysis in the continuous integration process can aid in approving potentially risky submissions.
@@
Benchmarking involves applying methods to data. Data types can be diverse and even a single dataset can be stored in multiple formats. A broad distinction can be made according to general vs. specific formats and open vs. proprietary formats. File formats specific to a programming language should be avoided since those formats may change over time, security risks can arise [51, 52], and they are not always interoperable. Likewise, proprietary formats should be avoided, since they may change and could require specific (proprietary) software to read. General language-agnostic formats often do not change over time and thus align to FAIR principles [7]. Generally, using established standards is recommended (e.g., SAM, BED, FASTQ, etc for genomics), since related methods are developed around these formats.
If raw input data consists of multiple file formats, an option could be to convert the data into a common format with the caveat that this ‘intermediate’ raw data could be either stored (increased storage usage, easier benchmark definition) or created ‘on the fly’ (increased compute). Furthermore, how raw input data are retrieved (e.g., automatic retrieval once and store locally, automatic retrieval every time, or manual retrieval and store locally) needs to be decided. These intermediate results could also be made FAIR.
Additionally, validation is recommended. Validations include content integrity (e.g., md5sum), adherence to a data specification (e.g., JSON Schema), and data semantics (e.g., SHACL for linked data).
Finally, metadata (automatic versus manual annotation) should not be neglected. Automatic generation of metadata might require predefined standards and an ontology to fully track how the data was generated [53].
@@ Conclusion
Benchmarking is a cornerstone practice in bioinformatics, impacting how methods are conceived, developed, published and chosen. Benchmarking is driven by the effort of people with the expertise to address biological and technical problems. This expertise can be leveraged by current practices of community benchmarking, so scientific advances are eased by joint efforts to catalog open problems, generate (or simulate) appropriate input data, develop methods to address them, and evaluate their performances. Community benchmarking has been proven successful and productive [54, 13, 14].
Here, we discussed the motivation and process of community benchmarking at several levels, from definition to execution, while highlighting strategies and technologies to potentially ease the technical and scientific challenges as well as how to organize contributions from communities.